An aqueous solution contains 30.0 g of a protein in 1.00 L. The osmotic pressure of the solution is 0.0167 atm at 27oC. What is the approximate molar mass of the protein?
1
Expert's answer
2020-12-14T07:18:13-0500
"pV=\frac{m}{M}RT\\(0.0167\cdot 101325)(10^{-3})=\frac{0.03}{M}(8.31)(27+273)\\M=44.2\frac{kg}{mol}"